cookiemonster:

Part A: Solve -vp + 30 < 45 for v.. Show your work. (4 points) Part B: Solve 3w – 6r = 30 for r. Show your work. (6 points)

cookiemonster:

@Vocaloid @Vocaloid @Ultrilliam @ThisGirlPretty

cookiemonster:

@Vocaloid please help

Vocaloid:

so on the left hand side you have -vp + 30 and you want v by itself how do we "get rid of" that + 30?

cookiemonster:

umm....-30 on both sides?

Vocaloid:

great, let's do that

Vocaloid:

-vp + 30 < 45 -vp < 15

Vocaloid:

how can we "get rid" of that "p"?

cookiemonster:

um divide?...

Vocaloid:

good

Vocaloid:

so now we have -v < 15/p now we just multiply both sides by -1 (making sure we flip the inequality around) giving is v > -15/p as our answer

Vocaloid:

now for 3w – 6r = 30 for r how can we "get rid of" that 3w on the left?

cookiemonster:

ummm -6r on both sides?

Vocaloid:

leave the -6r alone for now how can we "remove" the 3w?

cookiemonster:

3w subtract from both sides?

Vocaloid:

good, that gives us: – 6r = 30 - 3w how can we "get rid of" the -6 multiplied to r on the left hand side?

cookiemonster:

divide

Vocaloid:

ok, divide by what number?

cookiemonster:

-6r

Vocaloid:

just -6 since we want to leave r alone

cookiemonster:

oh right..

Vocaloid:

so we have: – 6r = 30 - 3w r = -(30-3w)/6

cookiemonster:

-5?

Vocaloid:

or more neatly:

Vocaloid:

r = (3w-30)/6

Vocaloid:

we can't simplify this further because we don't know what w is

Vocaloid:

so that's it

cookiemonster:

yay awesome thanks